OpenStudy (anonymous):

divide flowing polynomials (27x^4 - 18x^3 + 9x^2) ÷ 3x

OpenStudy (anonymous):

@amistre64 i got the answer on wolfram some how it doesn't fit

OpenStudy (anonymous):

zale help? i will give you medal if u will help

OpenStudy (zale101):

how many times does 3x go into 27x^4

OpenStudy (anonymous):

not sure :/

OpenStudy (zale101):

http://www.purplemath.com/modules/polydiv2.htm

OpenStudy (zale101):

how many times 3x goes into 27x^4 is obviosly 9x^2

zepdrix (zepdrix):

Remember how we did the problem earlier clay? :O We'll split up the fraction the same way we did before. \[\large \frac{27x^4-18x^3+9x^2}{3x} \qquad=\qquad \frac{27x^4}{3x}-\frac{18x^3}{3x}+\frac{9x^2}{3x}\]

zepdrix (zepdrix):

Zale was showing you how to do the first term. Divide the coefficients, 27/3 and also divide the x's. Remember that when you divide terms of similar bases, you `subtract` the exponents. \(\large \dfrac{x^4}{x^1}=x^{4-1}\)

OpenStudy (anonymous):

thnx zepdix :P enjoy dat pizza :P

zepdrix (zepdrix):

nom nom nom :3 You understand how to do the other terms? :o

OpenStudy (anonymous):

@zepdrix \[9x^3-6x^2+6 \]

OpenStudy (anonymous):

corrcet me if im wrong

zepdrix (zepdrix):

Hmm the first two terms look good. Let's look back over the last one though.\[\large \frac{9x^2}{3x} =\frac{9}{3}\cdot\frac{x^2}{x}\] Hmm 9/3 isn't 6 +_+

zepdrix (zepdrix):

You `subtract the exponents on x` but you don't subtract the coefficients. You divide those like you normally would.

OpenStudy (anonymous):

lol 3 then

zepdrix (zepdrix):

\[\large \frac{x^2}{x} \qquad=\qquad \frac{x^2}{x^1}\qquad=\qquad x^{2-1} \qquad=\qquad x^1 \qquad=\qquad x\]

OpenStudy (anonymous):

\[9x^3-6x^3+3 \]

OpenStudy (anonymous):

like that?

zepdrix (zepdrix):

No, you dropped the x for some reason. Didn't you see my last step....?

OpenStudy (anonymous):

type your answer

zepdrix (zepdrix):

\[\large \frac{9}{3}\cdot\frac{x^2}{x} \qquad=\qquad 3\cdot \frac{x^2}{x} \qquad=\qquad 3\cdot x\]

OpenStudy (anonymous):

look on screen

zepdrix (zepdrix):

\[\large [9][x][\;^3\;][-][6][x][\;^2\;][+][3][x]\]

OpenStudy (anonymous):

got it thats what i had but with out x